### v1

# The Hausdorff measure due to Davies and Rogers and its cardinal invariants

Abstract
Davies and Rogers constructed a Hausdorff measure satisfying the following property: every Borel subset of the space has measure either $\infty$ or 0. In this paper, we examine cardinal invariants of their measure.
1 Introduction
Hausdorff measures, which generalize the concept of the Lebesgue measure, is an important tool that can be useful in many fields. As for these, it is natural from a measure theoretic point of view to examine the claim that for any Hausdorff measure and any Borel set having infinite measure, there is a Borel subset of measure positive and finite inside it. In 1969, Davies and Rogers constructed a Hausdorff measure that is a counterexample to this. In this paper, we examine cardinal invariants of their measure.
The following is the standard definiton of cardinal invariants.
Definition 1.
  1. (1)
    We say $F\subseteq\omega^{\omega}$ is an unbounded family if $\neg(\exists g\in\omega^{\omega})(\forall f\in F)(f\leqslant^{*}g)$. Here $\leqslant^{*}$ is the almost domination order. Put $\mathfrak{b}=\min\{\lvert F\rvert:F\subseteq\omega^{\omega}\text{ unbounded family}\}$.
  2. (2)
    We say $F\subseteq\omega^{\omega}$ is a dominating family if $(\forall g\in\omega^{\omega})(\exists f\in F)(g\leqslant^{*}f)$. Put $\mathfrak{d}=\min\{\lvert F\rvert:F\subseteq\omega^{\omega}\text{ dominating family}\}$.
  3. (3)
    $\mathcal{N}$ and $\mathcal{M}$ denote the Lebesgue measure zero ideal and Baire first category ideal on $2^{\omega}$, respectively.
  4. (4)
    $\mathcal{E}$ denotes the $\sigma$-ideal on $2^{\omega}$ generated by closed Lebesgue null sets.
  5. (5)
    For an ideal $\mathcal{I}$ on a set $X$:
    1. (a)
      $\operatorname{add}(\mathcal{I})$ (the additivity of $\mathcal{I}$) is the smallest cardinality of a family $F$ of sets in $\mathcal{I}$ such that the union of $F$ is not in $\mathcal{I}$.
    2. (b)
      $\operatorname{cov}(\mathcal{I})$ (the covering number of $\mathcal{I}$) is the smallest cardinality of a family $F$ of sets in $\mathcal{I}$ such that the union of $F$ is equal to $X$.
    3. (c)
      $\operatorname{non}(\mathcal{I})$ (the uniformity of $\mathcal{I}$) is the smallest cardinality of a subset $A$ of $X$ such that $A$ does not belong to $\mathcal{I}$.
    4. (d)
      $\operatorname{cof}(\mathcal{I})$ (the cofinality of $\mathcal{I}$) is the smallest cardinality of a family $F$ of sets in $\mathcal{I}$ that satisfies the following condition: for every $A\in\mathcal{I}$, there is $B\in F$ such that $A\subseteq B$.
2 Review of Hausdorff measures
Definition 2.
We say $f\colon[0,\infty)\to[0,\infty)$ is a gauge function if it is nondecreasing and right-continuous and $f(0)=0$ holds.
Definition 3.
Let $(X,d)$ be a metric space, $f$ be a gauge function , $\delta>0$ be a real number. For $A\subseteq X$ and a sequence $\langle C_{n}:n\in\omega\rangle$} of subsets of $X$, we say this sequence is $\delta$-cover of $A$ if
$A\subseteq\bigcup_{n}C_{n}\text{ and }\mathrm{diam}(C_{n})\leqslant\delta\text{ for every $n$}.$
For a subset $A\subseteq X$, let
$\mathcal{H}^{f}_{X,\delta}(A)=\inf\left\{\sum_{n\in\omega}f(\mathrm{diam}(C_{n})):\text{${\langle C_{n}:n\in\omega\rangle}$ is a $\delta$-cover of $A$}\right\},$
which we call $\delta$-approximation of $f$-Hausdorff measure of $A$. Also let
$\mathcal{H}^{f}_{X}(A)=\sup_{\delta>0}\mathcal{H}^{f}_{X,\delta}(A),$
which we call $f$-Hausdorff measure of $A$.
If the context is clear, we omit the subscript $X$.
Remark 4.
Both $\delta$-approximation of $f$-Hausdorff measure and $f$-Hausdorff measure are outer measures.
Remark 5.
For every metric space $(X,d)$, every gauge function $f$ and every real number $\delta>0$, $\mathcal{H}^{f}(A)=0$ holds iff $\mathcal{H}^{f}_{\delta}(A)=0$ holds.
3 General results on additivity and cofinality
Let $(X,d)$ be a separable metric space and let $f$ be a gauge function. We call the following condition Assumption(*): there exist a constant $\alpha>1$ and a countable family $\mathcal{C}$ of subsets of $X$ such that for every subset $A\subseteq X$ of finite diameter and every $\varepsilon>0$, there is some $C\in\mathcal{C}$ with
$A\subseteq C\quad\text{and}\quad f(\mathrm{diam}(C))\leqslant\varepsilon+\alpha\cdot f(\mathrm{diam}(A)).$
Lemma 6.
Let $(X,d)$ be a separable metric space. Then there exists a countable family $\mathcal{C}$ of subsets of $X$ such that for every $A\subseteq X$ of finite diameter and every $\varepsilon>0$, there is some $C\in\mathcal{C}$ with
$A\subseteq C\text{ and }\mathrm{diam}(C)\leqslant\varepsilon+2\mathrm{diam}(A).$
Moreover, if $X$ is also an ultrametric space, then we can strengthen this to
$A\subseteq C\text{ and }\mathrm{diam}(C)\leqslant\varepsilon+\mathrm{diam}(A).$
Proof.
By separability, we can take a countable dense subset $D=\{a_{n}:n\in\omega\}$ of $X$. Let
$\mathcal{C}=\{B(a_{n},q):n\in\omega,\ q\in\mathbb{Q},\ q>0\}.$
We show that this $\mathcal{C}$ is the desired family.
Let $A\subseteq X$ have finite diameter and let $\varepsilon>0$ be given. The case $A=\varnothing$ is clear. Assume $A\neq\varnothing$ and take $b\in A$. By density of $D$, choose $n$ such that $d(a_{n},b)<\varepsilon/4$. Choose a positive rational $q$ such that
$\mathrm{diam}(A)+\varepsilon/4\leqslant q<\mathrm{diam}(A)+\varepsilon/2.$
We claim that $C=B(a_{n},q)$ is a desired member.
First, we show $A\subseteq C$. Let $a\in A$. Then
$d(a_{n},a)\leqslant d(a_{n},b)+d(b,a)<\varepsilon/4+\mathrm{diam}(A)\leqslant q,$
so $a\in C$.
Since $C$ is an open ball of radius $q$, the triangle inequality gives
$\mathrm{diam}(C)\leqslant 2q\leqslant\varepsilon+2\mathrm{diam}(A).$
If $X$ is an ultrametric space, then
$\mathrm{diam}(C)\leqslant q\leqslant\varepsilon+\mathrm{diam}(A).$
Corollary 7.
Each of the following pairs consisting of a separable metric space $(X,d)$ and a gauge function $f$ satisfies Assumption(*):
  1. (1)
    $f$ is a continuous doubling gauge function; that is, there exists $r>0$ such that
    $f(2x)<rf(x)\quad\text{for all }x\in[0,\infty).$
  2. (2)
    $(X,d)$ is an ultrametric space.
Proof.
(1) Take the countable family $\mathcal{C}$ from the lemma, and let $r>0$ witness the doubling condition. Let $A\subseteq X$ have finite diameter and let $\varepsilon>0$. Since $f$ is right-continuous, choose $\delta>0$ such that
$f(\mathrm{diam}(A)+\delta)-f(\mathrm{diam}(A))<\varepsilon/r.$
By the choice of $\mathcal{C}$, there exists $C\in\mathcal{C}$ with $A\subseteq C$ and
$\mathrm{diam}(C)\leqslant 2\delta+2\mathrm{diam}(A).$
Then
$f(\mathrm{diam}(C))\leqslant f(2\delta+2\mathrm{diam}(A))\leqslant rf(\delta+\mathrm{diam}(A))\leqslant\varepsilon+rf(\mathrm{diam}(A)).$
Hence Assumption(*) holds.
(2) Take the countable family $\mathcal{C}$ given by the lemma for ultrametric spaces. Let $A\subseteq X$ have finite diameter and let $\varepsilon>0$. Since $f$ is right-continuous, choose $\delta>0$ such that
$f(\mathrm{diam}(A)+\delta)-f(\mathrm{diam}(A))<\varepsilon.$
By the choice of $\mathcal{C}$, there exists $C\in\mathcal{C}$ with $A\subseteq C$ and
$\mathrm{diam}(C)\leqslant\delta+\mathrm{diam}(A).$
Then
$f(\mathrm{diam}(C))\leqslant f(\delta+\mathrm{diam}(A))\leqslant\varepsilon+f(\mathrm{diam}(A)).$
Hence Assumption(*) holds. ∎
Lemma 8.
Let $(X,d)$ be a separable metric space and let $f$ be a gauge function satisfying Assumption(*). Then for $A\subseteq X$, the following are equivalent: $A$ has $f$-Hausdorff measure 0 if and only if for every $\varepsilon>0$ there exists a sequence ${\langle C_{n}:n\in\omega\rangle}\subseteq\mathcal{C}$ such that
$A\subseteq\bigcup_{n\in\omega}C_{n}\quad\text{and}\quad\sum_{n\in\omega}f(\mathrm{diam}(C_{n}))<\varepsilon.$
Proof.
The “if” direction is clear. We prove the “only if” direction.
Assume $A$ has $f$-Hausdorff measure 0, and let $\varepsilon>0$. By definition, we can find ${\langle A_{n}:n\in\omega\rangle}\subseteq\mathcal{P}(X)$ such that
$A\subseteq\bigcup_{n\in\omega}A_{n}\quad\text{and}\quad\sum_{n}f(\mathrm{diam}(A_{n}))<\varepsilon/\alpha.$
Take a sequence of positive reals $\langle\eta_{n}:n\in\omega\rangle$} such that
$\sum_{n}\bigl(\eta_{n}+\alpha\cdot f(\mathrm{diam}(A_{n}))\bigr)\leqslant\varepsilon.$
For each $n$, choose $C_{n}\in\mathcal{C}$ such that $A_{n}\subseteq C_{n}$ and
$\mathrm{diam}(C_{n})\leqslant\eta_{n}+\alpha\cdot\mathrm{diam}(A_{n}).$
Then we obtain $\sum_{n}\mathrm{diam}(C_{n})\leqslant\varepsilon$. ∎
Proposition 9.
Let $(X,d)$ be a separable metric space and let $f$ be a gauge function. Assume Assumption(*) holds for these. Then $\mathcal{N}^{f}_{X}\preceq_{\mathrm{T}}\mathcal{N}$. In particular, under this assumption,
$\operatorname{add}(\mathcal{N})\leqslant\operatorname{add}(\mathcal{N}^{f}_{X})\quad\text{and}\quad\operatorname{cof}(\mathcal{N}^{f}_{X})\leqslant\operatorname{cof}(\mathcal{N}).$
Proof.
This is based on Theorem 534B of [2],
Let $\mathcal{S}=\prod_{n\in\omega}[\omega]^{\leqslant 2^{n}}$. By Bartoszyński’s characterization of the null ideal,
$(\mathcal{N},\mathcal{N},\subseteq)\equiv_{\mathrm{T}}(\omega^{\omega},\mathcal{S},\in^{*}),$
so it suffices to show $\mathcal{N}^{f}_{X}\leqslant_{\mathrm{T}}(\omega^{\omega},\mathcal{S},\in^{*})$. Take the countable family $\mathcal{C}$ from the assumption.
For each $n\in\omega$, let
$\mathcal{I}_{n}=\left\{I\subseteq\mathcal{C}\ \text{finite}:\ \sum_{C\in I}f(\mathrm{diam}(C))\leqslant 4^{-n}\right\}.$
Let $\langle I_{n,j}:j\in\omega\rangle$} be an enumeration of $\mathcal{I}_{n}$.
For $S\in\mathcal{S}$, define
$\psi(S)=\bigcap_{m\in\omega}\ \bigcup_{n\geqslant m}\ \bigcup_{i\in S(n)}\ \bigcup I_{n,i}.$
We show that $\psi(S)\in\mathcal{N}^{f}$.
$\mathcal{H}^{f}_{\infty}(\psi(S))$ $\leqslant\sum_{n\geqslant m}\ \sum_{i\in S(n)}4^{-n}=\sum_{n\geqslant m}4^{-n}\cdot 2^{n}=2^{-m}\to 0\ \ (m\to\infty),$
so this holds. Thus $\psi:\mathcal{S}\to\mathcal{N}^{f}_{X}$ is defined. Fix $A\in\mathcal{N}^{f}_{X}$. For each $n\in\omega$, we can take a sequence ${\langle C_{n,i}:i\in\omega\rangle}\subseteq\mathcal{C}$ such that
$A\subseteq\bigcup_{i}C_{n,i}\text{ and }\sum_{i}f(\mathrm{diam}(C_{n,i}))\leqslant 2^{-n}.$
Rearranging the double sequence $\langle C_{n,i}:n,i\in\omega\rangle$} into a single sequence, we obtain a sequence $\langle C_{i}:i\in\omega\rangle$} such that
$\sum_{i}f(\mathrm{diam}(C_{i}))\leqslant 1\text{ and }A\subseteq\bigcap_{m}\bigcup_{i\geqslant m}C_{i}.$
Choose a sequence of natural numbers $\langle k(n):n\in\omega\rangle$} satisfying
$\sum_{i=k(n)}^{\infty}f(\mathrm{diam}(C_{i}))\leqslant 4^{-n}\text{ for every }n.$
Define $\varphi(A)\in\omega^{\omega}$ by
$\{C_{i}:k(n)\leqslant i<k(n+1)\}=I_{n,\varphi(A)(n)}\text{ for each }n.$
Then $\varphi:\mathcal{N}^{f}_{X}\to\omega^{\omega}$ is defined and satisfies
$A\subseteq\bigcap_{m}\bigcup_{n\geqslant m}\bigcup I_{n,\varphi(A)(n)}.$
Therefore, for any $A\in\mathcal{N}^{f}$ and $S\in\mathcal{S}$, the condition $\varphi(A)\in^{*}S$ implies $A\subseteq\psi(S)$. This is what we needed to show. ∎
4 Review of the construction due to Davies and Rogers
In this section, we review the construction due to Davies and Rogers [1].
Lemma 10.
Let $n>0$ be an natural number. Then there is a finite graph $G$ that satisfies the following properties:
  1. (1)
    $G$ cannot be partitioned into $n$ independent subsets.
  2. (2)
    For every function $w\colon G\to\mathbb{R}_{\geqslant 0}$, there is an independent subset $H\subseteq G$ such that $\sum_{g\in H}w(g)\geqslant\frac{1}{4}\sum_{g\in G}w(g).$
  3. (3)
    There is a family $\langle H(x):x\in G\rangle$} of independent subsets such that
    1. (a)
      $x\in H(x)$ for every $x\in G$, and
    2. (b)
      $\lvert\{x\in G:a\in H(x)\}\rvert\geqslant\frac{1}{4}\lvert G\rvert$ for every $a\in G$.
Proof.
The proof is almost the same as in [1], but we have an additional condition (3). To prove it, we modify the original proof slightly and give a probabilistic proof.
Let $S^{n}=\{x\in\mathbb{R}^{n+1}:\|x\|=1\}$ be the $n$-sphere, and let $\sigma^{n}$ be the uniform measure on $S^{n}$. Set $\varepsilon_{n}=1/(2\sqrt{n})$.
For $x\in S^{n}$, define
$C(x)=\{y\in S^{n}:x\cdot y\geqslant\varepsilon_{n}\},$
and call it the cap centered at $x$.
For all sufficiently large $n$ and all $x\in S^{n}$, we have $\sigma^{n}(C(x))>\frac{1}{4}$ (see [1] for this computation). Refix such a $n$ larger than given $n$.
Let $m$ be a sufficiently large integer. Consider choosing $m$ points on $S^{n}$ independently at random with respect to $\sigma^{n}$. Fix a realization of these $m$ points, and let $G$ be the vertex set of a graph. We put an edge between $x,y\in G$ if and only if
$\|x-y\|\geqslant 2-\varepsilon_{n}^{2}.$
Then for any $x\in S^{n}$, the set $C(x)\cap G$ is an independent set.
We prove (1) of the lemma. Assume that $G$ is partitioned into $n$ subsets $G_{1},\dots,G_{n}$. For $i=1,\dots,n$, define
$H_{i}=\{h\in S^{n}:\text{there exists }g\in G_{i}\text{ with }\|h-g\|\leqslant\varepsilon_{n}\}.$
Each $H_{i}$ is a closed subset of $S^{n}$. Since $m$ is sufficiently large and the points are random, the sets $H_{i}$ cover $S^{n}$. Now apply the Lusternik–Schnirelmann–Borsuk theorem (if $S^{n}$ is covered by at most $n+1$ closed sets, then one of them contains an antipodal pair). Hence there exist $i\in\{1,\dots,n\}$ and $h\in S^{n}$ such that $\pm h\in H_{i}$. Thus we can find $g,g^{\prime}\in G_{i}$ with $\|h-g\|\leqslant\varepsilon_{n}$ and $\|h+g^{\prime}\|\leqslant\varepsilon_{n}$. It follows that $g$ and $g^{\prime}$ are adjacent, and (1) is proved.
Next we prove (2). We may assume that $w(g)>0$ for some $g\in G$, and also normalize so that $\sum_{g\in G}w(g)=1$. It suffices to find an independent set $H\subseteq G$ such that $\sum_{g\in H}w(g)\geqslant\frac{1}{4}$. Since $C(x)\cap G$ is independent, it is enough to show that there exists $x\in S^{n}$ with
$\sum_{g\in C(x)\cap G}w(g)\geqslant\frac{1}{4}.$
Let $c=\sigma^{n}(C(x))$ (this does not depend on $x$). Then
$\int_{S^{n}}\sum_{g\in C(x)\cap G}w(g)\,d\sigma^{n}(x)$ $=\sum_{g\in G}\int_{g\cdot x\geqslant\varepsilon_{n}}w(g)\,d\sigma^{n}(x)$
$=\sum_{g\in G}\left(w(g)\int_{g\cdot x\geqslant\varepsilon_{n}}d\sigma^{n}(x)\right)$
$=c>\frac{1}{4}.$
Therefore, for some $x\in S^{n}$ we have $\sum_{g\in C(x)\cap G}w(g)>\frac{1}{4}$, proving (2).
For (3), take $H(x)=C(x)\cap G$ and use $\sigma^{n}(C(x))>\frac{1}{4}$ together with a Monte-Carlo type argument. ∎
For each natural number $n$, fix one of the finite graphs $G$ obtained by substituting it into the Lemma 10 and let $G(n)$ be this graph.
We define a sequence $\langle M_{n},G_{n},N_{n}:n\in\omega\rangle$} as follows:
$M_{0}$ $=1$
$G_{n}$ $=G(M_{n})$
$N_{n}$ $=\lvert G_{n}\rvert$
$M_{n+1}$ $=2N_{n}\text{ (for $n\in\omega$)}$
Note that
$\lim_{n\to\infty}\frac{N_{0}\dots N_{n-1}}{M_{0}\dots M_{n}}=0\text{ and }\lim_{n\to\infty}\frac{N_{0}\dots N_{n}}{M_{0}\dots M_{n}}=\infty.$
The first one follows from $M_{n+1}=2N_{n}$ and the second one follows from $N_{n}\geqslant\frac{4}{3}M_{n}$, which is a consequence of (1) and (2) of Lemma 10.
Let $\Omega=\prod_{n\in\omega}G_{n}$.
Induce a metric $\rho$ into $\Omega$ as follows: For $x,y\in\Omega,x\neq y$, letting $n$ be the first index that $x$ and $y$ differ,
[…]
Define a gauge function $h$ as follows:
$h(0)=0,h(2^{-n})=\frac{1}{M_{0}\dots M_{n}}\text{ (for $n\in\omega$)}.$
For points other than these, define the value of $h$ by interpolating linearly.
Davies and Rogers proved the following two propositions.
Proposition 11.
$\mathcal{H}_{\Omega}^{h}(\Omega)=\infty$.
Proposition 12.
For every subset $A\subseteq\Omega$, we have either $\mathcal{H}_{\Omega}^{h}(A)=0$ or $\mathcal{H}_{\Omega}^{h}(A)=\infty$.
They used the following notion to prove above two propositions.
Let $S\subseteq\Omega$ be a subset and assume that $S$ is of the form:
$S=\prod_{i\in\omega}H_{i},$
and for some $j\in\omega$ we have
  • For every $i<j$, $H_{i}$ is a singleton,
  • $H_{j}\neq\varnothing$, and
  • For every $i>j$, we have $H_{i}=G_{i}$.
If $H_{j}$ is an independent subset of $G_{j}$, then we say $S$ is a standard set of rank $j$.
Note that standard sets of rank $j$ have diameter $2^{-j}$.
Lemma 13.
For every subset $S$ of $\Omega$, if $\lvert S\rvert\geqslant 2$, then there is a standard set containing $S$ and having same diameter.
5 $\mathsf{ZFC}$ results
Theorem 14.
$\mathcal{N}^{h}_{\Omega}$ is orthogonal to $\mathcal{E}$. In particular, $\operatorname{cov}(\mathcal{N}^{h}_{\Omega})\leqslant\operatorname{non}(\mathcal{E})$ and $\operatorname{cov}(\mathcal{E})\leqslant\operatorname{non}(\mathcal{N}^{h}_{\Omega})$ hold.
Proof.
For every $n\in\omega$, let $\mu_{n}$ be the uniform measure on $G_{n}$. Let $\mu$ be the product measure on $\Omega$ of $\langle\mu_{n}:n\in\omega\rangle$}.
For every $n\in\omega$ and $x\in G_{n}$, let $H(x)$ be one of the independent set $H$ found by (3) of Lemma 10 such that $x\in H$.
Define a Borel set $B$ as follows:
$B=\{(x,y)\in\Omega^{2}:(\exists^{\infty}n)\ y(n)\in H(x(n))\}.$
It is sufficient to prove that
  1. (1)
    For every $x\in\Omega$, $\mathcal{H}^{h}_{\Omega}(B_{x})=0$ holds.
  2. (2)
    For every $y\in\Omega$, $\mu(\Omega\smallsetminus B^{y})=0$ holds.
First, we observe that
$\mathcal{H}^{h}_{\Omega,1}(B_{x})$ $\leqslant\mathcal{H}^{h}_{\Omega,1}(\{(x,y)\in\Omega^{2}:\exists n\geqslant n_{0}\ y(n)\in H(x(n))\})$
$\leqslant\sum_{n\geqslant n_{0}}\mathcal{H}^{h}_{\Omega,1}(\{(x,y)\in\Omega^{2}:y(n)\in H(x(n))\})$
$\leqslant\sum_{n\geqslant n_{0}}\frac{N_{0}\dots N_{n-1}}{M_{0}\dots M_{n}}\leqslant\sum_{n\geqslant n_{0}}\frac{1}{M_{0}}\left(\frac{1}{2}\right)^{n}\to 0\text{ ($n_{0}\to\infty$)}.$
So we have (1).
Second, for $y\in\Omega$, we consider the set
$\Omega\smallsetminus B^{y}=\{x\in\Omega:\forall^{\infty}n\ y(n)\not\in H(x(n))\}.$
By independence, we have
$\mu(\Omega\smallsetminus B^{y})=\prod_{n\in\omega}\mu_{n}(\{a\in G_{n}:y(n)\not\in H(a)\})$
But we know that $\mu_{i}(\{a\in G_{i}:y(i)\not\in H(a)\})\leqslant\frac{3}{4}$ by (3) of Lemma 10. Thus we have (2). ∎
Therefore, we can draw the diagram as follows:
$\aleph_{1}$$\operatorname{add}(\mathcal{N})$$\operatorname{cov}(\mathcal{N})$$\operatorname{add}(\mathcal{M})$$\mathfrak{b}$$\operatorname{non}(\mathcal{M})$$\operatorname{cov}(\mathcal{M})$$\mathfrak{d}$$\operatorname{cof}(\mathcal{M})$$\operatorname{non}(\mathcal{N})$$\operatorname{cof}(\mathcal{N})$$\mathfrak{c}$$\operatorname{add}(\mathcal{N}^{h}_{\Omega})$$\operatorname{cof}(\mathcal{N}^{h}_{\Omega})$$\operatorname{cov}(\mathcal{N}^{h}_{\Omega})$$\operatorname{non}(\mathcal{N}^{h}_{\Omega})$$\operatorname{non}(\mathcal{E})$$\operatorname{cov}(\mathcal{E})$
6 Concictency Results
Zapletal [4] proved the following theorem.
Theorem 15.
Let $P_{\mathcal{N}^{h}_{\Omega}}$ be the idealized forcing of $\mathcal{N}^{h}_{\Omega}$. Then, $P_{\mathcal{N}^{h}_{\Omega}}$ is $\omega^{\omega}$-bounding and it adds no splitting reals.
Therefore, in the model obtained by countable support iteration of $P_{\mathcal{N}^{h}_{\Omega}}$ of length $\omega_{2}$, we have $\aleph_{1}=\mathfrak{d}<\operatorname{cov}(\mathcal{N}^{h}_{\Omega})=\mathfrak{c}=\aleph_{2}$. If we assume large cardinals, then $\mathfrak{r}=\aleph_{1}$ holds in the model obtained by the above iteration using Theorem 6.3.8 of [4].
7 Consistency of $\operatorname{cov}(\mathcal{E})<\operatorname{non}(\mathcal{N}^{h}_{\Omega})$
Employing the following theorem due to Klausner and Mejia, we prove the consistency of $\operatorname{cov}(\mathcal{E})<\operatorname{non}(\mathcal{N}^{h}_{\Omega})$.
Theorem 16 ([3]).
Let $c,H,d\in\omega^{\omega}$ and assume $c>H\geqslant^{*}1$, $d\geqslant 2$ and $\limsup_{n\to\infty}\frac{\log_{d(n)}H(n)}{d(n)}=\infty$. Then there is a proper forcing poset $Q_{c,H}^{d}$ such that the following conditions hold:
  1. (1)
    $Q_{c,H}^{d}$ adds a slalom in $\mathcal{S}(c,H)$ that catches infinitely often every real in $\prod c$ in the ground model. In particular, the countable support iteration of $Q_{c,H}^{d}$ of length $\omega_{2}$ increases $\mathfrak{v}^{\exists}_{c,H}$.
  2. (2)
    Assume additionally the following conditions:
    1. (a)
      $a,e\in\omega^{\omega}$, $a>0$, $e$ goes to infinity.
    2. (b)
      $\prod_{k<n}a(k)\leqslant d(n)$ and $\prod_{k<n}c^{\triangledown H}(k)\leqslant e(n)$ for all but finitely many $n$.
    3. (c)
      $\lim_{k\to\infty}\min\{\frac{c^{\triangledown H}(k)}{e(k)},\frac{a(k)}{d(k)}\}=0$.
    Then, $Q_{c,H}^{d}$ is $(a,e)$-bounding, that is it forces every $x\in\prod{a}$ there is a slalom in $\mathcal{S}(a,e)$ in the ground model that catches $x$ eventually. Combining this result and Shelah’s preservation theorem, we have that the countable support iteration of $Q_{c,H}^{d}$ of length $\omega_{2}$ forces $\mathfrak{c}^{\forall}_{a,e}=\aleph_{1}$.
Lemma 17.
If $a,e\in\omega^{\omega}$ with $0<e\leqslant a$ satisfies $\prod_{n\in\omega}\frac{e(n)}{a(n)}=0$, then $\operatorname{cov}(\mathcal{E})\leqslant\mathfrak{c}^{\forall}_{a,e}$.
Proof.
Consider the uniform measure $\mu_{n}$ on $a(n)$ for every $n$ and $\mu$ the product measure on $\prod a$. Since $\operatorname{cov}(\mathcal{E})=\operatorname{cov}(\mathcal{E}_{\mu})$ holds, it suffices to show that there is Tukey morphism $(\varphi,\psi)\colon(\prod a,\mathcal{E}_{\mu},\in)\to(\prod a,\mathcal{S}(a,e),\in^{*})$, that is
there is $\varphi\colon\prod a\to\prod a$ and $\psi\colon\mathcal{S}(a,e)\to\mathcal{E}_{\mu}$ such that
$\hskip 28.45274pt\text{$\varphi(x)\in^{*}S$ implies $x\in\psi(S)$ for every $x\in\prod a$ and $S\in\mathcal{S}(a,e)$}.$
Letting $\varphi(x)=x$ and $\psi(S)=\{x\in\prod a:(\forall^{\infty}n)\ x(n)\in S(n)\}$ for $x\in\prod a$ and $S\in\mathcal{S}(a,e)$ suffices. Note that $\mu(\psi(S))=0$ for every $S$ by using the assumption of this lemma. ∎
In the following lemma, the sequence $\langle M_{n},N_{n}:n\in\omega\rangle$} is the one fixed when $\mathcal{N}^{h}_{\Omega}$ is defined.
Lemma 18.
Let $c,H\in\omega^{\omega}$ and $\langle I_{n}:n\in\omega\rangle$} be an interval partition. Assume $c(n)=\prod_{k\in I_{n}}N_{k}$ and $M_{\min{I_{n+1}}}>2^{n}H(n)N_{0}\dots N_{\min I_{n}-1}$ for every $n$. Then there is a Tukey morphism from $(\Omega,\mathcal{N}^{h}_{\Omega},\in)$ to $(\prod c,\mathcal{S}(c,H),\in^{\infty})$. In particular $\operatorname{cov}(\mathcal{N}^{h}_{\Omega})\leqslant\mathfrak{c}^{\exists}_{c,H}$ and $\mathfrak{v}^{\exists}_{c,H}\leqslant\operatorname{non}(\mathcal{N}^{h}_{\Omega})$ hold.
Proof.
Write $I_{n}=[i_{n},i_{n+1})$ for $n\in\omega$. Define $\varphi\colon\Omega\to\prod c$ and $\psi\colon\mathcal{S}(c,H)\to\mathcal{N}^{h}_{\Omega}$ as follows:
$\varphi(x)$ $=x,$
$\psi(S)$ $=\{x\in\Omega:(\exists^{\infty}n)\ (x\upharpoonright I_{n}\in S(n))\}.$
We must show that $\mathcal{H}^{h}_{\Omega}(\psi(S))=0$ for every $S$. But by the assumption, we have
$\frac{H(n)N_{0}\dots N_{i_{n}-1}}{M_{0}\dots M_{i_{n+1}}}\leqslant 2^{-n}.$
Thus, we have
$\sum_{n\in\omega}\frac{H(n)N_{0}\dots N_{i_{n}-1}}{M_{0}\dots M_{i_{n+1}}}<\infty.$
On the other hand, we have $\mathcal{H}^{h}_{\Omega,\infty}(\{x:x\upharpoonright I_{n}\in S(n)\})\leqslant\frac{H(n)N_{0}\dots N_{i_{n}-1}}{M_{0}\dots M_{i_{n+1}}}$. Therefore, we have $\mathcal{H}^{h}_{\Omega}(\psi(S))=0$. ∎
Theorem 19.
It is consistent that $\operatorname{cov}(\mathcal{E})<\operatorname{non}(\mathcal{N}^{h}_{\Omega})$.
Proof.
By induction, we define $c,H,d,a,e\in\omega^{\omega}$ and an interval partition $\langle I_{n}:n\in\omega\rangle$} as follows:
  1. (1)
    Define $d(0)=2,e(0)=1$.
  2. (2)
    Define $H(n)=(d(n))^{nd(n)}$.
  3. (3)
    Pick $\max I_{n}$ so that $M_{\min{I_{n+1}}}>2^{n}H(n)N_{0}\dots N_{\min I_{n}-1}$ and $\prod_{k\in I_{n}}N_{k}>H(n)$.
  4. (4)
    Define $c(n)=\prod_{k\in I_{n}}N_{k}$.
  5. (5)
    Define $a(n)=2e(n)$.
  6. (6)
    Define $e(n+1)=2\prod_{k\leqslant n}2^{c(k)}$.
  7. (7)
    Define $d(n+1)=\prod_{k\leqslant n}a(k)$.
Then, we have all assumptions of Theorem 16, Lemma 17 and Lemma 18. Therefore, the countable support iteration of $Q_{c,H}^{d}$ of length $\omega_{2}$ forces
$\aleph_{1}=\operatorname{cov}(\mathcal{E})\ =\mathfrak{c}^{\forall}_{a,e}<\aleph_{2}=\mathfrak{v}^{\exists}_{c,H}=\operatorname{non}(\mathcal{N}^{h}_{\Omega})=\mathfrak{c}.\qed$
Lemma 18 has another corollary:
Corollary 20.
In Mathias model, $\operatorname{cov}(\mathcal{N}^{h}_{\Omega})=\aleph_{1}$ holds. In particular, it is consistent that $\aleph_{1}=\operatorname{cov}(\mathcal{N}^{h}_{\Omega})<\mathfrak{b}=\mathfrak{s}=\mathfrak{c}=\aleph_{2}$.
Proof.
By the Laver property, in Mathias model, $\mathfrak{c}^{\forall}_{c,H}$ is small. In particular $\mathfrak{c}^{\exists}_{c,H}$ is also small. Thus this corollary follows from Lemma 18. ∎
Theorem 21.
Every finite support iteration of $\sigma$-centered forcings does not add $\mathcal{N}^{h}_{\Omega}$-random real.
Proof.
This proof is based on the standard proof that $\sigma$-center forcings do not add the (usual) random real.
We define a Polish relational system $R$ such that $\mathfrak{b}(R)=\operatorname{cov}(\mathcal{N}^{h}_{\Omega})$ and every $\sigma$-centered forcing has goodness for the relation $R$.
First we define
$D=\{f\colon\omega\to\operatorname{StdSets}:\sum_{n}h(\mathrm{diam}(f(n)))\leqslant 1\},$
Here, $\operatorname{StdSets}$ is the set of all standard sets. For $f\in D$ and $x\in\Omega$, we define
$f\sqsubseteq^{\mathcal{N}^{h}_{\Omega}}_{n}x\iff\forall k\geqslant n\ x\not\in f(k).$
Also, we define $\sqsubseteq^{\mathcal{N}^{h}_{\Omega}}=\bigcup_{n\in\omega}\sqsubseteq^{\mathcal{N}^{h}_{\Omega}}_{n}$. Therefore, putting
$A_{f}=\{x\in\Omega:\exists^{\infty}n\ x\in f(n)\},$
we have the equivalece
$f\sqsubseteq^{\mathcal{N}^{h}_{\Omega}}x\iff x\not\in A_{f}.$
Let $R=(D,\Omega,\sqsubseteq^{\mathcal{N}^{h}_{\Omega}})$. It is clear that $R$ is a Polish relational system and $\mathfrak{b}(R)=\operatorname{cov}(\mathcal{N}^{h}_{\Omega})$.
So in the remaining part, we show every $\sigma$-centered forcing has goodness for $R$. Let $P$ be a $\sigma$-centered forcing and $P=\bigcup_{m\in\omega}P_{m}$, where each $P_{m}$ is a centered subset of $P$. Let $\dot{x}$ be a $P$-name such that $P\Vdash\dot{x}\in\Omega$. Let $N$ be a countable elementary substructure of a sufficient large model containing $\dot{x}$ and $\langle P_{m}:m\in\omega\rangle$}. Also let $f\in D$ be a $\sqsubseteq^{\mathcal{N}^{h}_{\Omega}}$-dominating element over $N$. We have to show $P\Vdash\dot{x}\in A_{f}$. Let $p_{0}\in P$ and take $m$ such that $p_{0}\in P_{m}$. For each $n\in\omega$, we put
$S^{n}=\{s\in\prod_{i<n}G_{i}:\forall p\in P_{m}\ \exists q\leqslant p\ q\Vdash\dot{x}\upharpoonright n=s\}.$
We claim here that each $S_{n}$ is nonempty. If $S_{n}$ were empty, then for every $s\in\prod_{i<n}G_{i}$, there would exist $p_{s}\in P_{m}$ forcing $\dot{x}\upharpoonright n\neq s$. Taking a common extension of $\{p_{s}:s\in\prod_{i<n}G_{i}\}$ would give a contradiction.
Therefore, $\bigcup_{n}S_{n}$ is an infinite subtree of a compact tree $\bigcup_{n}\prod_{i<n}G_{i}$. Thus, by König’s lemma, there is $y\in\Omega$ such that $y\upharpoonright n\in S_{n}$ for every $n$. This $y$ can be taken so that $y\in N$. Therefore, it holds that $y\in A_{f}$. Fix a natural number $k$ such that $y\in f(k)$. Since $f(k)$ is open, we can take $\ell$ such that $y\upharpoonright\ell]\subseteq f(k)$. Since $y\upharpoonright\ell\in S^{\ell}$, we can take $q\leqslant p$ such that $\dot{x}\upharpoonright\ell=y\upharpoonright\ell$. Therefore $q\Vdash\dot{x}\in f(k)$. Since $k$ is an arbitrary number such that $y\in f(k)$ and there exist infinitely many such $k$, we have $q\Vdash\dot{x}\in A_{f}$. Since the condition $p$ were arbitrary, we have $P\Vdash\dot{x}\in A_{f}$. ∎
8 Questions
The following questions remain.
Question 22.
  1. (1)
    Do we have $\mathsf{Con}(\operatorname{add}(\mathcal{N})<\operatorname{add}(\mathcal{N}^{h}_{\Omega}))$ and $\mathsf{Con}(\operatorname{cof}(\mathcal{N}^{h}_{\Omega})<\operatorname{cof}(\mathcal{N}))$?
  2. (2)
    Do we have $\operatorname{add}(\mathcal{N}^{h}_{\Omega})\leqslant\mathfrak{b}$ and $\mathfrak{d}\leqslant\operatorname{cof}(\mathcal{N}^{h}_{\Omega})$?
  3. (3)
    Do we have $\mathsf{Con}(\operatorname{add}(\mathcal{N}^{h}_{\Omega})<\operatorname{cov}(\mathcal{N}^{h}_{\Omega}))$ and $\mathsf{Con}(\operatorname{non}(\mathcal{N}^{h}_{\Omega})<\operatorname{cof}(\mathcal{N}^{h}_{\Omega}))$?
  4. (4)
    Do we have $\mathsf{Con}(\operatorname{non}(\mathcal{N}^{h}_{\Omega})<\operatorname{cov}(\mathcal{N}^{h}_{\Omega}))$?

### v2

@@
       $\operatorname{cof}(\mathcal{I})$ (the cofinality of $\mathcal{I}$) is the smallest cardinality of a family $F$ of sets in $\mathcal{I}$ that satisfies the following condition: for every $A\in\mathcal{I}$, there is $B\in F$ such that $A\subseteq B$.
+The cardinal invariants defined below are so-called localization cardinals and anti-localization cardinals.
+Definition 2.
+  1. (1)
+    For $c\in(\omega+1)^{\omega}$ and $h\in\omega^{\omega}$, define $\prod c=\prod_{n\in\omega}c(n)$ and $S(c,h)=\prod_{n\in\omega}[c(n)]^{\leqslant h(n)}$.
+  2. (2)
+    For $x\in\prod c$ and $\varphi\in S(c,h)$, define $x\in^{*}\varphi$ iff $(\forall^{\infty}n)(x(n)\in\varphi(n))$ and define $x\in^{\infty}\varphi$ iff $(\exists^{\infty}n)(x(n)\in\varphi(n))$.
+  3. (3)
+    $\mathfrak{c}^{\forall}_{c,h}=\min\{\lvert S\rvert:S\subseteq S(c,h),(\forall x\in\prod c)(\exists\varphi\in S)(\forall^{\infty}n)(x(n)\in\varphi(n))\}$.
+  4. (4)
+    $\mathfrak{c}^{\exists}_{c,h}=\min\{\lvert S\rvert:S\subseteq S(c,h),(\forall x\in\prod c)(\exists\varphi\in S)(\exists^{\infty}n)(x(n)\in\varphi(n))\}$.
+  5. (5)
+    $\mathfrak{v}^{\forall}_{c,h}=\min\{\lvert X\rvert:X\subseteq\prod c,(\forall\varphi\in S(c,h))(\exists x\in X)(\exists^{\infty}n)(x(n)\not\in\varphi(n))\}$.
+  6. (6)
+    $\mathfrak{v}^{\exists}_{c,h}=\min\{\lvert X\rvert:X\subseteq\prod c,(\forall\varphi\in S(c,h))(\exists x\in X)(\forall^{\infty}n)(x(n)\not\in\varphi(n))\}$.
 2 Review of Hausdorff measures
@@
 which we call $f$-Hausdorff measure of $A$.
+We put $\mathcal{N}^{f}_{X}=\{A\subseteq X:\mathcal{H}^{f}_{X}(A)=0\}$.
 If the context is clear, we omit the subscript $X$.
@@
 Theorem 15.
-$\mathcal{N}^{h}_{\Omega}$ is orthogonal to $\mathcal{E}$.
-In particular, $\operatorname{cov}(\mathcal{N}^{h}_{\Omega})\leqslant\operatorname{non}(\mathcal{E})$ and $\operatorname{cov}(\mathcal{E})\leqslant\operatorname{non}(\mathcal{N}^{h}_{\Omega})$ hold.
+$\operatorname{cov}(\mathcal{N}^{h}_{\Omega})\leqslant\operatorname{non}(\mathcal{E})$ and $\operatorname{cov}(\mathcal{E})\leqslant\operatorname{non}(\mathcal{N}^{h}_{\Omega})$ hold.
 Proof.
@@
     2. (b)
-      $\prod_{k<n}a(k)\leqslant d(n)$ and $\prod_{k<n}c^{\triangledown H}(k)\leqslant e(n)$ for all but finitely many $n$.
+      $\prod_{k<n}a(k)\leqslant d(n)$ and $\prod_{k<n}c^{\triangledown H}(k)\leqslant e(n)$ for all but finitely many $n$. Here $c^{\triangledown H}(k)$ means $\lvert[c(k)]^{\leqslant h(k)}\rvert$.
     3. (c)
@@
 ∎
+Corollary 23.
+It is consistent that $\operatorname{cov}(\mathcal{N}^{h}_{\Omega})<\operatorname{add}(\mathcal{M})$.
+Also i is consistent that $\operatorname{cof}(\mathcal{M})<\operatorname{non}(\mathcal{N}^{h}_{\Omega})$.
+Proof.
+Consider finite support iterations of Hechler forcing of length $\omega_{2}$ and length $\omega_{2}+\omega_{1}$, respectively. ∎
 8 Questions
 The following questions remain.
-Question 22.
+Question 24.
   1. (1)